• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 07:22:00 on October19,2009

    Entry number 295644

    keyword=Photon background rate

    Background rate have been higher and more unstable since the mid owl shift, the number of HV trip of the detector kept increasing during the week-end. MCC people checked the orbit and pathlength but did not find anything in particular. I guess people from the injector will check if something changed on their end when the experts are in.
    Figure attached is background ( laser off ) with a gain of 8 and HV at 1500V which roughly corresponds to the usual running conditions of 1800 V and attenuation by a factor of 3

    FIGURE 1